OpenStudy (anonymous):

could anyone help me with the half angle formulas for tan. i have tan(5pi)/4 i can get everything plugged into the formula but i am having trouble simplify @precal

OpenStudy (ajprincess):

Do u want to find the value of tan(5pi)/4?

OpenStudy (ajprincess):

@bfinn4

OpenStudy (anonymous):

yes the exact value

OpenStudy (anonymous):

i am having trouble simplify all the radical stuff..and i dont have a way to type a radical into this web site so i dont know how to show you my work?

OpenStudy (anonymous):

oops i typed that woring i want the value of tan(5pi)/8

OpenStudy (anonymous):

or tan [((5pi)/4)/2]

OpenStudy (ajprincess):

`\(sqrt{}\)` u can use this to type radicals.

OpenStudy (ajprincess):

\[\tan\theta=\frac{2tan(\frac{\theta}{2})}{1-tan^2(\frac{\theta}{2})}\]

OpenStudy (ajprincess):

\[\frac{\theta}{2}=\frac{5\pi}{8}\] \[\theta=\frac{5\pi}{4}\] plugging this in the formula gives u \[\tan\frac{5\pi}{4}=\frac{2tan\frac{5\pi}{8}}{1-tan^2\frac{5\pi}{8}}\] \[\pi=180\] so \[\frac{5\pi}{4}=225\] do u knw the value of tan225?

OpenStudy (anonymous):

1

OpenStudy (ajprincess):

yup good:)

OpenStudy (ajprincess):

so \[\tan\frac{5\pi}{4}=1\]. \[1=\frac{2tan\frac{5\pi}{8}}{1-tan^2\frac{5\pi}{8}}\] Let us assume \[x=\tan\frac{5\pi}{8}\] \[1=\frac{2x}{1-x^2}\] \[1-x^2=2x\] \[x^2+2x-1=0\] Can u find x?

OpenStudy (ajprincess):

@bfinn4

OpenStudy (anonymous):

i have to go to class but yes i can find x ...i will do it right after geology.

OpenStudy (ajprincess):

k 5n:)

OpenStudy (precal):

@ajprincess she did a great job helping you